OpenStudy (anonymous):

HELP ME PLEASE!!!!!!! A company produces accessories for smart phones and tablets. The profit on each smart phone case is $2, and the profit on each tablet case is $3. The company made a profit of $1,200 on the cases last month. The equation 2x + 3y = 1,200 represents the company's profit from cases last month, where x is the number of smart phone cases sold and y is the number of tablet cases sold.   1. Change the equation into slope-intercept form. Identify the slope and y-intercept of the equation. Be sure to show all of your work.

OpenStudy (anonymous):

Describe how you would graph this line using the slope-intercept method. Be sure to write in complete sentences. Write the equation in function notation. Explain what the graph of the function represents. Be sure to use complete sentences. Graph the function using one of the following two options below. On the graph, make sure to label the intercepts. You may graph your equation by hand on a piece of paper and scan your work. You may graph your equation using graphic technology that can be found in the Course Information area. Suppose in the next month, the total profit on smart phone cases and tablet cases is $1,500. The profit amounts are the same, $2 for smart phone case and $3 for the tablet case. In a paragraph of at least three sentences, explain how the graphs of the functions for the two months are the same and how they are different. Be sure to use complete sentences. Below is a graph that represents the total profits for a third month. Write the equation of the line that represents this graph. Show your work or explain how you determined the equations.

OpenStudy (anonymous):

@shinalcantara

OpenStudy (shinalcantara):

Slope intercept form is defined by the equation, y = mx + b where m is the slope of the line and b is the y-intercept.. Following the form for the slope-intercept, then we'll have: 2x+3y = 1200 subtracting 2x to each side of the equation we'll have 2x + 3y - 2x = 1200 - 2x 3y = 1200 -2x for y to have a coefficient of 1 as indicated by the form we divide the whole equation with 3 making it y = 1200/3 - 2/3x or \[y = -\frac{ 2 }{ 3 }x + 400\] Knowing that b=400 then we'll have the point for the y-intercept as (0,400). In finding the x-intercept we will use the slope formula since we already have the slope and a point and we know that an x-intercept is having a value of y=0. \[m = \frac{ y-y_1 }{ x-x_1 }\] \[-\frac{ 2 }{ 3 } = \frac{ 0 - 400 }{ x-0 }\] \[\frac{ -2 }{ 3 } = \frac{ -400 }{ x }\] \[x = \frac{ -400(3) }{ -2 }\] \[x = 600\] so we'll have x-intercept as (600,0)

OpenStudy (shinalcantara):



OpenStudy (anonymous):

Thanks!